# Subvarieties of pointed Abelian $\ell$-groups

### v1

@@ 4 Characterization of varieties generated by a single finitely generated totally ordered $\ell$-group
    For the other inclusion assume that some equation $\alpha(\vec{x})=0$ does not hold in ${\boldsymbol{\mathit{R}}}_{a}$. Thus there is $\vec{r}\in\mathbb{R}^{m}$, where $m$ is the number of variables in $\alpha$, such that $\alpha(\vec{r})\neq 0$. By Lemma 4.2 the function $\vec{x}\mapsto\alpha(\vec{x})$ is continuous. Since the set $\{x\in\mathbb{R}\ \mid x\neq 0\}$ is open in $\mathbb{R}$, the set $O:=\{\vec{x}\in\mathbb{R}^{m}\mid\alpha(\vec{x})\neq 0\}$ is open in $\mathbb{R}^{m}$. We know that $\vec{r}\in O$ thus the set $O$ is nonempty. Since $A$ is a dense subset of $\mathbb{R}$, by [26, Theorem 19.5] $A^{m}$ is a dense subset of $\mathbb{R}^{m}$ and therefore $A^{m}\cap O\neq\emptyset$. Thus there exist $\vec{a}\in A^{m}\cap O$ such that $\alpha(\vec{a})\neq 0$. This shows that the equation $\alpha(\vec{x})=0$ is not valid in ${\boldsymbol{\mathit{A}}}_{a}$. Since $\alpha(\vec{x})=0$ was an arbitrary equation we derive that ${\boldsymbol{\mathit{R}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})$. Thus $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{a})$.
@@
    For the other inclusion assume that some equation $\alpha(\langle{\vec{y},\vec{x}}\rangle)=0$ does not hold in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}$. Let $m$ be the number of variables in $\alpha$. There exist $\langle{\vec{s},\vec{r}}\rangle\in A^{m}\times\mathbb{R}^{m}$ such that $\alpha(\langle{\vec{s},\vec{r}}\rangle)\neq 0$ in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}$. By Lemma 4.2 the function $\vec{x}\rightarrow\alpha(\vec{s},\vec{x})$ is continuous and thus the set $O:=\{\vec{x}\in{\boldsymbol{\mathit{R}}}^{m}\mid\alpha(\langle{\vec{s},\vec{x}}\rangle)\neq 0\}$ is open in ${\boldsymbol{\mathit{R}}}^{m}$ and contains $\vec{r}$. Therefore, there is $\epsilon\in\mathbb{R}$ such that $V:=\{\vec{x}\mid|\vec{x}-\vec{r}|<m\cdot\epsilon\}\subseteq O$. We fix $n\in I$ such that $n>\frac{\sqrt{m}}{2\epsilon}$. Let us consider $\ell$-group ${{\boldsymbol{\mathit{Z}}}[\frac{1}{n}]}_{1}$. We have $\epsilon>\frac{\sqrt{m}}{2n}$ and by [10, Section 5, Chapter 4] it follows that ${{\boldsymbol{\mathit{Z}}}[\frac{1}{n}]}_{1}\cap V\neq\emptyset$. Since $V\subseteq O$ the equation $\alpha(\vec{s},\vec{x})=0$ is not valid in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{{\boldsymbol{\mathit{Z}}}[\frac{1}{n}]}_{1}$. By Lemma 4.1 we get ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{{\boldsymbol{\mathit{Z}}}[\frac{1}{n}]}_{1}\cong{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}$, which proves ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}\in\mathbf{HSP}(\{{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}\mid n\in I\})$.
@@
    This can be proved similarly as the previous point.
@@
This shows that $\varphi$ preserves the ordering. In similar way one can show that $\varphi^{-1}$ preserves the ordering as well. Thus $\varphi$ is indeed an isomorphism. ∎
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered pointed Abelian $\ell$-group and $a\neq 0$. We call the unique maximal convex ideal $I$ of ${\boldsymbol{\mathit{A}}}_{a}$ such that $a\notin I$ a p-radical, and we denote it by $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$. Furthermore, we say that ${\boldsymbol{\mathit{A}}}_{a}$ is p-simple if it has a trivial p-radical, $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})=\{0\}$.
For any Abelian pointed $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$, its radical, $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$, is an Abelian (non-pointed) $\ell$-group. Determining a canonical point for $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ isn’t immediately obvious. However, when $\boldsymbol{\mathit{A}}$} is finitely generated, Lemma 3.3 provides the decomposition ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\overrightarrow{\times}(\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))_{b}$ for a unique $b\in\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$. This means that for any finitely generated $\ell$-group, we can meaningfully consider its radical as a pointed Abelian $\ell$-group. We will define its point, $\mathtt{f}^{\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})}$, such that the decomposition ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ holds. The preceding decomposition makes it possible to treat $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ as a pointed Abelian $\ell$-group. Henceforth, we adopt the convention that when we do so, its point is the one uniquely determined by this decomposition.
Let us note that here the definition of p-simple and p-radical does not coincide with the classical universal algebraic meaning of an $\ell$-group being simple and the notion of the radical of $\ell$-group. When we say that a pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$ is p-simple or $\boldsymbol{\mathit{C}}$} is a p-radical of ${\boldsymbol{\mathit{A}}}_{a}$, we are saying something much weaker than an Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} being simple or that $\boldsymbol{\mathit{C}}$} is a radical of $\boldsymbol{\mathit{A}}$}. One can show that the notion of p-radical and radical coincide (as well the notions of simple and p-simple) in the case of $\ell$-groups with a strong unit.
@@
This definition gives us the following useful property: Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered strongly pointed Abelian $\ell$-group and let ${\boldsymbol{\mathit{B}}}_{a}$ be its totally ordered pointed Abelian $\ell$-subgroup and $a\neq 0$. It can be shown that $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{B}}}_{a})$. Since ${\boldsymbol{\mathit{B}}}_{a}$ is strongly pointed, the quotient ${\boldsymbol{\mathit{B}}}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{B}}}_{a})$ has to be a simple $\ell$-group and hence, according to Hölder’s Theorem (see Theorem 2.8), ${\boldsymbol{\mathit{B}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{B}}}_{a})$ is isomorphic to a pointed $\ell$-subgroup of the real numbers. Since any pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$ with $a\neq 0$ has a unique convex strongly pointed $\ell$-subgroup, we can state the following definition.
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered pointed Abelian $\ell$-group and ${\boldsymbol{\mathit{B}}}_{a}$ be its convex strongly pointed $\ell$-subgroup. We define rank of ${\boldsymbol{\mathit{A}}}_{a}$ as
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered p-simple pointed Abelian $\ell$-group and $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n\in\mathbb{Z}\setminus\{0\}$. Then ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{n}$.
@@
Follows from the definitions of rank and p-simple $\ell$-group. ∎
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered non-p-simple Abelian $\ell$-group with rank $n$. Then $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$.
@@
By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Since $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n$, by Lemma 3.2 and Lemma 4.1 we obtain ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$. We first show ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$.
@@
Since ${\boldsymbol{\mathit{A}}}_{a}$ is finitely generated, $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ has a strong unit and thus Lemma 3.5 gives us ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}(\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$ and thus by Lemma 3.9 we obtain ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$. Finally, ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$ and thus ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$.
@@
We have to show the other inclusion, i.e. that ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. By Lemma 4.4 we have ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{kn}\in\mathbf{I}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ for each $k\in\mathbb{Z}$. Thus by Lemma 4.3 we obtain ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1},{\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{-1}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. Also, by Lemma 2.11 we have ${\boldsymbol{\mathit{R}}}_{0}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})$ and thus by Lemma 3.9 and Lemma 4.1 we have ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{0}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$.
@@
Since $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ is totally ordered, by Lemma 2.11 we get $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$, thus by [2, Theorem 5.6] we have
$\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{1}).$
@@
${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{-1})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{0})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}).$

### v2

@@ 3 The lexicographic decompositions
    For the other inclusion assume that some equation $\alpha(\vec{x})=0$ does not hold in ${\boldsymbol{\mathit{R}}}_{a}$. Thus there is $\vec{r}\in\mathbb{R}^{m}$, where $m$ is the number of variables in $\alpha$, such that $\alpha(\vec{r})\neq 0$. By Lemma 4.2 the function $\vec{x}\mapsto\alpha(\vec{x})$ is continuous. Since the set $\{x\in\mathbb{R}\ \mid x\neq 0\}$ is open in $\mathbb{R}$, the set $O:=\{\vec{x}\in\mathbb{R}^{m}\mid\alpha(\vec{x})\neq 0\}$ is open in $\mathbb{R}^{m}$. We know that $\vec{r}\in O$ thus the set $O$ is nonempty. Since $A$ is a dense subset of $\mathbb{R}$, by [30, Theorem 19.5] $A^{m}$ is a dense subset of $\mathbb{R}^{m}$ and therefore $A^{m}\cap O\neq\emptyset$. Thus there exists $\vec{a}\in A^{m}\cap O$ such that $\alpha(\vec{a})\neq 0$. This shows that the equation $\alpha(\vec{x})=0$ is not valid in ${\boldsymbol{\mathit{A}}}_{a}$. Since $\alpha(\vec{x})=0$ was an arbitrary equation we derive that ${\boldsymbol{\mathit{R}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})$. Thus $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{a})$.
@@
    For the other inclusion assume that some equation $\alpha(\langle{\vec{y},\vec{x}}\rangle)=0$ does not hold in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}$. Let $m$ be the number of variables in $\alpha$. There exist $\langle{\vec{s},\vec{r}}\rangle\in A^{m}\times\mathbb{R}^{m}$ such that $\alpha(\langle{\vec{s},\vec{r}}\rangle)\neq 0$ in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}$. By Lemma 4.2 the function $\vec{x}\rightarrow\alpha(\vec{s},\vec{x})$ is continuous and thus the set $O:=\{\vec{x}\in{\boldsymbol{\mathit{R}}}^{m}\mid\alpha(\langle{\vec{s},\vec{x}}\rangle)\neq 0\}$ is open in ${\boldsymbol{\mathit{R}}}^{m}$ and contains $\vec{r}$. Therefore, there is $0<\epsilon\in\mathbb{R}$ such that $V:=\{\vec{x}\mid|\vec{x}-\vec{r}|<m\cdot\epsilon\}\subseteq O$. We fix $n\in I$ such that $n>\frac{\sqrt{m}}{2\epsilon}$. Let us denote $\boldsymbol{\mathit{B}}$} the pointed $\ell$-subgroup of $\boldsymbol{\mathit{R}}$} generated by $\frac{1}{n}$. We have $\epsilon>\frac{\sqrt{m}}{2n}$ and by [12, Section 5, Chapter 4] it follows that ${\boldsymbol{\mathit{B}}}_{1}^{m}\cap V\neq\emptyset$. Since $V\subseteq O$ the equation $\alpha(\vec{s},\vec{x})=0$ is not valid in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{1}$. Since ${\boldsymbol{\mathit{B}}}_{1}\cong{\boldsymbol{\mathit{Z}}}_{n}$ by Lemma 4.1 we get ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{1}\cong{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}$, which proves ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}\in\mathbf{HSP}(\{{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}\mid n\in I\})$.
@@
    This can be proved similarly as the previous point. ∎
@@
This shows that $\varphi$ preserves the ordering. In a similar way one can show that $\varphi^{-1}$ preserves the ordering as well. Thus $\varphi$ is indeed an isomorphism. ∎
@@
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a pointed Abelian $\ell$-group with $b\neq 0$. A convex $\ell$-subgroup of $\boldsymbol{\mathit{A}}$} that is maximal with respect to not containing $b$ is called a value of $b$. If ${\boldsymbol{\mathit{A}}}_{b}$ is totally ordered, the value of $b$ is unique and we denote it by $\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})$. Furthermore, we say that ${\boldsymbol{\mathit{A}}}_{b}$ is p-simple if $\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})=\{0\}$.
For any pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{b}$, $\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})$ is not a pointed Abelian $\ell$-group. To treat $\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})$ itself as an algebra in $p\mathbb{AL}$, we must assign it a designated element. When $\boldsymbol{\mathit{A}}$} is finitely generated, Lemma 3.3 applies to establish ${\boldsymbol{\mathit{A}}}\cong{\boldsymbol{\mathit{A}}}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})$. Under this isomorphism, the original designated element $b\in A$ must map to a specific pair. The first coordinate is its projection $b+\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})$ in the quotient, while the second coordinate is a uniquely determined element $c\in\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})$. We use this point $c$ as the designated element for $\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})$. Henceforth, we adopt the convention that whenever $\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})$ is treated as a pointed Abelian $\ell$-group, its designated element $\mathtt{f}^{\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})}$ is precisely this element $c$, yielding:
${\boldsymbol{\mathit{A}}}_{b}\cong({\boldsymbol{\mathit{A}}}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{b}))_{b+\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})}\overrightarrow{\times}(\mathbf{val}({\boldsymbol{\mathit{A}}}_{b}))_{c}.$
@@
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a totally ordered pointed Abelian $\ell$-group with $b\neq 0$ and let ${\boldsymbol{\mathit{B}}}_{b}$ be the totally ordered strongly pointed convex $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{b}$. It can be shown that $\mathbf{val}({\boldsymbol{\mathit{A}}}_{b})=\mathbf{val}({\boldsymbol{\mathit{B}}}_{b})$. Since ${\boldsymbol{\mathit{B}}}_{b}$ is strongly pointed, $\mathbf{val}({\boldsymbol{\mathit{B}}}_{b})$ is a maximal convex $\ell$-subgroup of $\boldsymbol{\mathit{B}}$} and hence the quotient ${\boldsymbol{\mathit{B}}}/\mathbf{val}({\boldsymbol{\mathit{B}}}_{b})$ must be a simple Abelian $\ell$-group. Therefore, according to Hölder’s Theorem (see Theorem 2.9), ${\boldsymbol{\mathit{B}}}_{b}/\mathbf{val}({\boldsymbol{\mathit{B}}}_{b})$ is isomorphic to a pointed $\ell$-subgroup of the real numbers. Since any pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{b}$ with $b\neq 0$ has a unique convex strongly pointed $\ell$-subgroup, we can state the following definition.
@@
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a totally ordered pointed Abelian $\ell$-group with $b\neq 0$ and ${\boldsymbol{\mathit{B}}}_{b}$ be its strongly pointed convex $\ell$-subgroup. We define the rank of ${\boldsymbol{\mathit{A}}}_{b}$ as
@@
Moreover, for any totally ordered 0-pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{0}$ we set $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{0})=0$.
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered strongly pointed Abelian $\ell$-group and $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n\in\mathbb{Z}\setminus\{0\}$. Moreover, assume ${\boldsymbol{\mathit{A}}}_{a}$ is p-simple. Then ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{n}$.
@@
Follows from the definitions of rank and p-simple Abelian $\ell$-group. ∎
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered pointed Abelian $\ell$-group which is not p-simple and has rank $n\in\mathbb{Z}\setminus\{0\}$. Then $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$.
@@
By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Since $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n$, by Lemma 3.2 and Lemma 4.1 we obtain ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\cong{\boldsymbol{\mathit{Z}}}_{n}$ and thus ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\cong{\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$. We first show ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a}))$.
@@
Since ${\boldsymbol{\mathit{A}}}_{a}$ is finitely generated and totally ordered, $\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$ has a strong unit and thus Lemma 3.5 gives us ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}(\mathbf{val}({\boldsymbol{\mathit{A}}}_{a}))$ and thus by Lemma 3.9 we obtain ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a}))$. Finally, we observe ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a}))$ and thus ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a}))$.
@@
We have to show the other inclusion, i.e. that ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. By Lemma 4.4 we have ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{kn}\in\mathbf{I}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ for each $k\in\mathbb{Z}$. Thus by Lemma 4.3 we obtain ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1},{\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{-1}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. Also, by Lemma 2.13 we have ${\boldsymbol{\mathit{R}}}_{0}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})$ and thus by Lemma 3.9 and Lemma 4.1 we have ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{0}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$.
@@
Since $\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$ is totally ordered, by Lemma 2.13 and [3, Theorem 5.6] we have
$\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{1}).$
@@
${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{-1})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{0})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}).$